Use FW analysis to compare between machines A and B if the MARR is 10%

Machine A Machine B
First Cost, $ -20,000 -30,000
Annual Cost, $ year -9000 -7000
Salvage Value, $ 4000 6000
Life 3 6

The Future Worth of machine A

The Future Worth of machine B

Which machine should be Selected

A.

$-68,960

B.

Machine B

C.

$-203,272

D.

$-101,156

E.

Machine A

F.

$-122,168

G.

$-103,245
Answer #1

Future Worth of A = -20000(F/P, 10%, 6) – 20000(F/P, 10%, 3) – 9000(F/A, 10%, 6) + 4000(F/P, 10%,

3) + 4000

= -20000*1.7716 – 20000*1.3310 - 9000*7.7156 + 4000*1.3310 + 4000

= -122,168

Future worth of B = -30000(F/P, 10%, 6) – 7000(F/A, 10%, 6) + 6000

= -30000*1.7716 – 7000*7.7156 + 6000

= - 101,156

Select B because it has a greater present worth at - 101,156